OpenStudy (anonymous):

a negative point charge of 2.0 * 10^-6 C experiences a force of 0.060 N from another source charge. What is the strength of the electric field, and how does its direction compare with the force? A. 1.2 * 10^3 N/C, same B. 1.5 * 10^4 N/C , opposite C. 3.0 * 10^4 N/C , opposite D. 3.0 * 10^4 N/C , same

OpenStudy (michele_laino):

here we have to apply this formula: \[\Large F = QE\] or: \[\Large E = \frac{F}{Q} = \frac{{6 \times {{10}^{ - 2}}}}{{2 \times {{10}^{ - 6}}}}\]

OpenStudy (michele_laino):

furthermore, since Q is positive, then vectors F and E have the same orientation

OpenStudy (anonymous):

30,000? ohh so it will be choice D?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yay!! thank you:)

OpenStudy (michele_laino):

:)